OpenStudy (anonymous):

Divide the quantity ten times ten to the seventeenth power divided by the quantity one hundred times ten to the thirteenth power. Express the answer in scientific notation. one point zero zero times ten to the third power one point zero zero times ten to the thirtieth power zero point one zero times ten to the fourth power zero point one zero times ten to the thirtieth power

OpenStudy (anonymous):

PLease help me i really really need it

OpenStudy (anonymous):

\[\frac{10^{17}}{100\times 10^{13}}=\frac{10^{17}}{10^{15}}=10^{17-15}=10^2=100\]

OpenStudy (anonymous):

since this is not one of your possible answers, i am going to assume that i misinterpreted the question that is how i read it

OpenStudy (anonymous):

wait so whats the answer i really need to get this done

OpenStudy (anonymous):

it has to be one of these one point zero zero times ten to the third power one point zero zero times ten to the thirtieth power

OpenStudy (campbell_st):

aren't you missing something \[\frac{10^1 \times 10^17}{ 10^2 \times 10^13} = \] just use the index law for multiplcation, add the powers and the index law for division, subtract the powers.